# What Color Is It? A Text-Interference Multimodal Hallucination Benchmark

### v1

@@ IV-B1 Recognition Accuracy
Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question, and $A$ be the answer provided by $M$. We define $A$ as ”correct” if and only if:
@@
where $\text{Text}(I)$ extracts the text content from $I$, and $\text{VisualColor}(T)$ denotes the visual color of the text $T$.
@@ IV-B2 Hallucination 1 — Hallucination Caused by Text Information Interference
Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question about text color, and $A$ be the answer (color) provided by $M$. The function $\text{Text}(I)$ extracts the text string in $I$. The function $\text{VisualColor}(\text{Text}(I))$ is the true visual color of the text in $I$. The set $\mathcal{S}(\text{Text}(I))$ is the **semantic color set** (i.e., colors commonly associated with the text content).
The model $M$ exhibits a ”Type I Hallucination” if and only if its answer $A$ satisfies the following two conditions:
@@ IV-B3 Hallucination 2 — Hallucination Caused by Question Color Interference
Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question, and $A$ be the color answer provided by $M$. The function $\text{Text}(I)$ extracts the text string in $I$. The set $\mathcal{S}(Q)$ is the **semantic color set** indicated by the content of the question $Q$.
The model $M$ exhibits a ”Type II Hallucination” if and only if its answer $A$ satisfies:

### v2

@@ IV-B1 Recognition Accuracy
Let $M$ be a multimodal model, II be an input image, $Q$ be a question, and AA be the answer provided by $M$. We define AA as ”correct” if and only if:
@@
where $\text{Text}(I)$ extracts the text content from II, and $\text{VisualColor}(T)$ denotes the visual color of the text $T$.
@@ IV-B2 Hallucination 1 — Hallucination Caused by Text Information Interference
Let $M$ be a multimodal model, II be an input image, $Q$ be a question about text color, and AA be the answer (color) provided by $M$. The function $\text{Text}(I)$ extracts the text string in II. The function $\text{VisualColor}(\text{Text}(I))$ is the true visual color of the text in II. The set $\mathcal{S}(\text{Text}(I))$ is the **semantic color set** (i.e., colors commonly associated with the text content).
The model $M$ exhibits a ”Type I Hallucination” if and only if its answer AA satisfies the following two conditions:
@@ IV-B3 Hallucination 2 — Hallucination Caused by Question Color Interference
Let $M$ be a multimodal model, II be an input image, $Q$ be a question, and AA be the color answer provided by $M$. The function $\text{Text}(I)$ extracts the text string in II. The set $\mathcal{S}(Q)$ is the **semantic color set** indicated by the content of the question $Q$.
The model $M$ exhibits a ”Type II Hallucination” if and only if its answer AA satisfies: